How do you solve for Surface Area of a square pyramid but not including the what touches the ground and the volume of the pyramid? The question just says that it is 230 meters across on all sides and has a height of 140 meters.
Answer by fractalier(6550)

Okay, can you picture the height going from the center of the bottom to the very top of the pyramid? If you draw a line from that point, along the base, you can see a right triangle that is 115 m at its base and 140 m tall, right?
The slant height of the pyramid's face can be figured using the Pythagorean Theorem...
Slant Height, call it c,
c^2 = a^2 + b^2 = 115^2 + 140^2
so that
c = 181.2 m
Now that slant height is the height of each triangular face, whose base is 230 m. The area of a triangle is
A = (1/2)bh = (1/2)(230)(181.2) = 20838 m^2
but there are four such faces so that the total surface area is
4A, or
83352 m^2